A vehicle travelling at a constant speed of 60 km/hr rounds a curve of radius 100 m. Find its acceleration?

A. 1.777 m/s2
B. 2.777 m/s2
C. 3.777 m/s2
D. 4.777 m/s2
E. 5. 777 m/s2

Kindly, tell me what should be its answer and why, please.

Thank you.
1
Expert's answer
2020-08-20T09:48:06-0400

The centrepetal acceleration

"a=\frac{v^2}{R}=\frac{(60\times \frac{1000\:\rm m}{3600\:\rm s})^2}{100\:\rm m}=2.777\:\rm m/s^2"

Answer: B